# Using iterated local alignment to aggregate trajectory data into a traffic flow map

### v2

@@ 3 Local alignment of road segments for flow map aggregation
Our goal of producing a minimal flow map relies on resolving the crucial problem of how to aggregate similar, but not exactly overlapping, road segments. Many solutions have been offered, such as edge bundling (Zhou et al., 2013) and rasterisation (Wood et al., 2010, Morgan and Lovelace, 2021). Edge bundling consists of clustering trajectory linestrings and replacing all cluster members with a single representative linestring. These (and subsequent) authors conclude that it performs poorly when applied to noisy trajectories at the road segment level, and remains mostly suited to coarser aggregations, such as desire lines. Rasterisation relies on converting the vectorial flow map into a raster matrix, and aggregating the flows within the same raster pixel neighbourhood. Whilst this is indeed able to improve flow aggregations, it depends highly on the raster pixel neighbourhood size, and the rasterisation of the vectorial flow map leads to a loss of spatial resolution. We propose an alternative aggregation which does not lose resolution. Due to the complexity of this aggregation, it is divided into several algorithms, so that after the application of each algorithm, we progress further towards a minimal flow map.
We develop our novel algorithms within the R statistical analysis environment, to take advantage of its integrated access to advanced statistical and geospatial analysis methods. Whilst R is not a bona fide GIS (Geographical Informations System), its geospatial functionalities conform to the OGC standards (OGC, 2010) via the package sf (Pebesma, 2018), and is a viable option for research in transport geospatial data analysis (Necula, 2015, Lovelace et al., 2019).
3.1 Node snapping with hierarchical clustering
@@
Since we are searching for points which are close to together, then this is well-suited to statistical clustering. There are many statistical clustering algorithms available, and we focus on hierarchical clustering (Gordon, 1999). A naive implementation where we consider all boundary points of all flow linestrings in a 1-pass complete linkage clustering is computationally intensive for any reasonable number of routes (Müllner, 2013). To resolve this computational bottleneck, we approximate the 1-pass complete linkage clustering by a nested 2-pass clustering. We begin with an efficient single linkage clustering of the boundary points of all linestrings in the R package fastcluster (Müllner, 2013). Since single linkage can result in chain-like clusters, we compute a subsequent complete linkage clustering to break these potential chains. In this nested approach, the complete linkage distance matrix is calculated only within each single linkage cluster, and so we are less likely to reach computational limits.
Algorithm 2 is a description of ST_SNAPNODE. The inputs are the $n_{\mathcal{F}}$ traffic flow linestrings =$\mathcal{F}=\{{\bm{f}}_{1},\dots,{\bm{f}}_{n_{\mathcal{F}}}\}$ and the snap tolerance $\varepsilon_{S}$. In Step 1, we extract the boundary points of all flow linestrings. In Steps 2–3, we compute a single linkage clustering on all boundary points, and cut the dendrogram at height $\varepsilon_{S}$, resulting in $C^{\prime}$ clusters. In Steps 4–7, within each of these single linkage clusters, we compute a complete linkage clustering, and cut the dendrogram at height $\varepsilon_{S}$. This divides each single linkage cluster into $C^{\prime\prime}$ clusters, where all cluster members are at most $\varepsilon_{S}$ distance from each other. In Steps 8–11, we compute the weighted spatial centroid in each of the $C^{\prime\prime}$ complete linkage clusters, weighted by the corresponding traffic flow values. In Step 12, we renumber the cluster labels from the above nested clusterings to approximate a 1-pass complete linkage clustering into $C^{*}$ clusters. In Steps 13–16, within each of these $C^{*}$ clusters, for all points of the corresponding flow linestrings which are closer than $\varepsilon_{S}$ distance to the centroid, we snap them to the centroid. In Steps 17–18, we collate and sort all the $n_{\mathcal{F}^{*}}$ snapped linestrings =$\mathcal{F}^{*}=\{{\bm{f}}_{1}^{*},\dots,{\bm{f}}_{n_{\mathcal{F}^{*}}}^{*}\}$. In comparison to the unsnapped linestrings $\mathcal{F}$, there are fewer snapped linestrings $\mathcal{F}^{*}$ which have higher flow values, are more interconnected, and share more exactly overlapping sub-segments.
@@
3:Extract boundary points $B(\mathcal{F}):=\{\operatorname{Start}({\bm{e}}_{1}),\operatorname{End}({\bm{e}}_{1}),\dots,\operatorname{Start}({\bm{e}}_{n_{\mathcal{F}}}),\operatorname{End}({\bm{e}}_{n_{\mathcal{F}}})\}$
@@
16:     Extract corresponding flow linestrings $\mathcal{F}_{i}^{*}:=\{{\bm{f}}_{1}^{*},\dots,{\bm{f}}_{n_{B^{*}}}^{*}\}$ from $\mathcal{F}$
17:     Snap points of $\{{\bm{f}}_{1}^{*},\dots,{\bm{f}}_{n_{B^{*}}}^{*}\}$ within $\varepsilon_{S}$ distance of cluster centroid ${\bm{b}}^{*}_{i}$ to ${\bm{b}}^{*}_{i}$
@@
19:Collate snapped linestrings $\{\mathcal{F}_{1}^{*},\dots\mathcal{F}_{C^{\prime*}}^{*}\}$ into :=$\mathcal{F}^{*}:=\{{\bm{f}}_{1}^{*},\dots,{\bm{f}}_{n_{\mathcal{F}^{*}}}^{*}\}$
@@
An illustration of Algorithm 2 is given in Figure 5. In Figure 5(a) is the flow map before node snapping, where we observe that the boundary points (black solid circles) of some of the purple linestrings are closer to each than the snapping tolerance $\varepsilon_{S}=4$ m. Figure 5(b) shows an update flow map after node snapping, and applying ST_OVERLINE_PLANR. There are fewer thin purple road segments, since by snapping their intersection nodes, they share more exact sub-segments, and these shared sub-segments are merged during the flow aggregation to result in wide orange road segments.
Refer to caption(a) Refer to caption(b)
Figure 5: Flow maps for unsnapped and node snapped segments, with tolerance $\varepsilon_{S}=4~{}\mathrm{m}$. (a) Unsnapped. (b) Node snapped. Colour (purple to orange) and width of road segments is proportional to traffic flow.
3.2 Node splitting to add missing intersection nodes
Node snapping is focused near the boundary points of the flow linestrings. So it does not address misalignments far from the boundary points. Moreover there remain some linestrings which do indeed intersect but whose intersection is not correctly computed by ST_OVERLINE_PLANR. The solution to both of these problems is the explicit addition of the missing intersection nodes to these linestrings.
ST_SPLITNODE is the collation of a couple of algorithms which explicitly add the nodes at the intersections in the interior of linestrings, and splits these linestrings into simple linestring segments at these added nodes. This gives the name to this method. The inputs to Algorithm 3 are the linestrings $\mathcal{F}$ and the node splitting $S$. For computational stability and efficiency, our method draws on two existing algorithms: in Steps 1–3, to_spatial_subdivision ($S=$ ‘subdivision’) in the sfnetworks package (van der Meer et al., 2023), and in Steps 4–5, geos_unary_union ($S=$ ‘unary’) in the geos package (Dunnington and Pebesma, 2023). The first option tends to find fewer intersection nodes, but this can be an advantage for our proposed line blending (to be introduced in the next subsection) since many small linestrings with similar flow values do not provide a clear prioritisation of this blending. It is also similar to the intersection computation performed by ST_OVERLINE_PLANR. The second option finds more missing intersection nodes, and leads to more comprehensive flow aggregation. We require both types of node splitting to compute a minimal flow map.
@@ 3.2 Node splitting to add missing intersection nodes
6:else if $S==$ ‘unary’ then
@@
In Figure 6(a) is the flow map without any node splitting or node snapping. This map has missing intersection nodes, and intersections nodes which are close to each other. In Figure 6(b) is the flow map after node splitting ($S=$ ‘unary’) to add the missing intersections nodes, and then node snapping (with snap tolerance $\varepsilon_{S}=4$ m). The result is that there are fewer road segments with higher flow values. Due to the combined action of ST_SPLITNODE and ST_SNAPNODE, Figure 6(b) is an improvement over Figures 5(a–b) and 6(a).
Refer to caption(a) Refer to caption(b)
Figure 6: Flow maps for unsplit/unsnapped and node split/node snapped segments, with tolerance $\varepsilon_{S}=4~{}\mathrm{m}$. (a) Unsplit and unsnapped. (b) Node split and node snapped. Colour (purple to orange) and width of road segments is proportional to traffic flow.
@@ 3.3 Line blending to align similar linestrings with local reference
So far we have focused on improving the alignment of linestrings induced by resolving inconsistencies at their intersections. We now focus on aligning linestrings more generally. For this, we require a comparison relationship to establish an order of alignment of nearby linestrings. For two linestrings ${\bm{f}}_{1}$ and ${\bm{f}}_{2}$ from a flow map $\mathcal{F}$, we define that ${\bm{f}}_{2}$ is a candidate to be aligned to the reference ${\bm{f}}_{1}$ at threshold $\varepsilon\geq 0$ if
@@
This relation is insensitive to any local complexities in ${\bm{f}}_{2}$ as long as they are all contained within the buffer zone around ${\bm{f}}_{1}$. The buffer zone we employ has flat edges, e.g. for the R package sf, this corresponds to ST_BUFFER(endCapStyle="FLAT"), so the buffer zone ends at the boundary points of ${\bm{f}}_{1}$. This avoids erroneously considering neighbouring segments of ${\bm{f}}_{1}$, which are connected to ${\bm{f}}_{1}$ as a part of a longer sequence of road segments, to be candidates to be aligned to ${\bm{f}}_{1}$. The condition on the flow values means that we place a higher priority on linestrings with higher flows. Since we can define a local reference linestring, a global road network graph is no longer required to align the candidate linestrings.
Let ${\bm{f}}_{\rm ref}$ be a reference linestring from $\mathcal{F}$. The set of $m$ linestrings from \$\mathcal{F}\backslash\{{\bm{f}}_{\rm ref}\}$ which satisfy Equation (1) are the candidate linestrings $\mathcal{F}_{\rm cand}=\{{\bm{f}}_{{\rm cand},1},\dots,{\bm{f}}_{{\rm cand},n_{\mathcal{F}_{\rm cand}}}\}$, with the convention that if $n_{\mathcal{F}_{\rm cand}}=0$ then $\mathcal{F}_{\rm cand}$ is the empty set. We call our approach ‘line blending’, since we will blend the candidate linestrings onto the reference linestring. The inputs in Algorithm 4 are the reference linestring ${\bm{f}}_{\rm ref}$, the set of $m$ candidate linestrings $\mathcal{F}_{\rm cand}$, and the blend tolerance $\varepsilon$. The output are the modified reference and $m$ modified candidate flow linestrings, all with added interior points for accurate calculation of exactly equal linestring segments. In Step 1, we initialise a local network graph $\mathcal{N}^{*}$ with the reference linestring. In Steps 2–3, we extract all points of the candidate linestrings, and use the network_blend function in the sfnetworks package (which we denote as ST_NETWORK_BLEND) to blend efficiently these points into the reference linestring (van der Meer et al., 2023). This network blending requires a blend threshold, which we set to be $\varepsilon$. ST_NETWORK_BLEND projects the candidate linestrings onto the reference linestring, and explicitly adds them to the network, thereby creating new edges in $\mathcal{N}^{*}$. The result is a local network graph with more, shorter edges and with nodes at the projected candidate points, and whose union is the reference linestring. In Steps 4–5, we extract the $m$ blended candidate linestrings with these added interior points by applying the shortest path search between the start and end point of each blended candidate linestring along the network graph $\mathcal{N}^{*}$. Step 6 is the equivalent for the blended reference linestring. Step 7 involves collating the blended candidate linestrings.
@@
2:Output: $({\bm{f}}^{*}_{\rm ref},\mathcal{F}^{*}_{\rm cand})$ blended reference and $m$ candidate flow linestrings
@@
8:${\bm{f}}^{*}_{{\rm ref}}$ := shortest path from $\operatorname{Start}({\bm{f}}_{{\rm ref}})$ to $\operatorname{End}({\bm{f}}_{{\rm ref}})$ along network $\mathcal{N}^{*}$
9:Collate $\mathcal{F}^{*}_{\rm cand}:=\{{\bm{f}}_{{\rm cand},1}^{*},\dots,{\bm{f}}_{{\rm cand},n_{\mathcal{F}_{\rm cand}}}^{*}\}$
@@
The result from ${\tt ST\_LINEBLEND}$ is the modified reference linestring and the projected candidate linestrings with added interior points. These added interior points resolve a limitation of the flow aggregation of ST_OVERLINE_PLANR (Morgan and Lovelace, 2021). If there are many candidate linestrings, then this may lead to many sub-segments in the projected linestrings, each with their own flow values. So we assign the weighted mean flow, weighted by the sub-segment length, to all sub-segments. This single flow value takes into account the contribution of each candidate linestring to the flow along the reference linestring. Moreover, we take the rounded value of this weighted mean flow to expedite the aggregation computations and to reduce visual clutter of the flow map.
@@
In Figure 7(a), we illustrate this line blending with the linestrings $ABC$ with flow $f_{ABC}=7$ in blue, and $AD$ with flow $f_{AD}=2$ in orange. As these two linestrings do not share a common sub-segment, so flow aggregation does not modify them. Since $AD\subset{\tt ST\_BUFFER}(ABC,4~{}\mathrm{m})$ (the pale blue rectangle), and $f_{AD}<f_{ABC}$, then $ABC$ is the reference ${\bm{f}}_{\rm ref}$ and $AD$ is the candidate linestring ${\bm{f}}_{\rm cand}$, according to Equation (1). Line blending involves projecting $AD$ to $ABC$, so $D$ is projected to $D^{\prime}$ (which lies exactly on the reference linestring), and $B$ is added explicitly to this projected linestring. We also add $D^{\prime}$ to the reference linestring. The reference linestring becomes ${\bm{f}}^{*}_{\rm ref}=(7,ABD^{\prime}C)$, and the candidate ${\bm{f}}^{*}_{\rm cand}=(2,ABD^{\prime})$. Now ${\bm{f}}^{*}_{\rm ref},{\bm{f}}^{*}_{\rm cand}$ share the sub-segment $ABD^{\prime}$ exactly, and ST_OVERLINE_PLANR gives the aggregated flows as $(9,ABD^{\prime})$ and $(2,D^{\prime}C)$. We take the rounded value of the weighted mean of these two flow linestrings. The result of ST_LINEBLEND with ST_OVERLINE_PLANR is a single linestring $(9,ABD^{\prime}C)$, as shown in Figure 7(b).
Refer to caption(a) Refer to caption(b)
@@
Figure 7: Line blending removes small misalignments and leads to correct flow aggregation. (a) Before line blending and flow aggregation. Reference linestring $(7,ABC)$ in blue, candidate linestring $(2,AD)$ in orange. (b) After blending candidate into reference linestring, with tolerance $\varepsilon=4$ m, and flow aggregation. Blended linestring is $(9,ABD^{\prime}C)$.
@@
More complex situations arise when other linestrings touch the candidate linestring, but are not candidates themselves for blending. Since line blending projects the candidates to the reference linestring, then we have to also project these other linestrings to avoid leaving a gap in the updated flow map. This procedure ST_SNAP_CAND_TOUCH is outlined in Algorithm 5. The inputs are the reference linestring ${\bm{f}}_{\rm ref}$, the $n_{\mathcal{F}_{\rm cand}}$ candidate linestrings $\mathcal{F}_{\rm cand}$, the $n_{\mathcal{F}_{\rm ct}}$ candidate-touching linestrings $\mathcal{F}_{\rm ct}$, and the snap tolerance $\varepsilon_{S}$. In Steps 1–8, we iterate over each candidate-touching linestring. In Steps 2–3, we compute the intersection points between the candidate-touching linestring and the boundary points of the candidate linestrings, and the respective distances. In Steps 4–7, if this intersection point is within $\varepsilon_{S}$ distance to the boundary points of ${\bm{f}}_{\rm ref}$, then we snap the candidate-touching linestring to the closest boundary point. In Step 8, if the intersection point is not within $\varepsilon_{S}$ distance, then we snap it to the nearest point on ${\bm{f}}_{\rm ref}$. These snapping operations are similar to those in Steps 15–16 in ST_SNAPNODE in Algorithm 2, and ensure that we maintain connectivity between $\mathcal{F}_{\rm ct}$ and ${\bm{f}}_{\rm ref}$. In Step 9, we collate these snapped linestrings.
@@
Figure 8(a) is an illustration of ST_SNAP_CAND_TOUCH with the reference $ABC$ with flow $f_{ABC}=7$ (blue), the candidate linestring $AD$ with flow $f_{AD}=2$ (orange), and the blending buffer zone with blend tolerance $\varepsilon=4$ m (light blue). The candidate-touching linestring ${\bm{f}}_{\rm ct}=(5,DE)$ (purple) touches the candidate $AD$ at $D$. The orange line is entirely within the pale blue buffer zone around the blue reference line, whereas the purple line extends outside of the buffer zone and so is not a candidate for blending to the reference linestring. If we apply ST_LINEBLEND to blend the candidate linestring, then $D$ is projected to $D^{\prime}$ to lie on the reference linestring, i.e. ${\bm{f}}_{\rm cand}^{*}=(2,ABD^{\prime})$. If we apply ST_SNAP_CAND_TOUCH to the candidate-touching linestring, then as the intersection point $D$ is less than $\varepsilon_{S}=4$ m from the boundary points of ${\bm{f}}_{\rm ref}$, it is projected to the boundary point $C$, i.e. ${\bm{f}}^{*}_{\rm ct}=(5,CE)$. Thus line blending, candidate-touching snapping and flow aggregation yields that ${\bm{f}}^{*}_{\rm ct}=(5,CE)$ remains connected to ${\bm{f}}_{\rm ref}^{*}=(9,ABD^{\prime}C)$.
Refer to caption(a) Refer to caption(b)

### v5

@@ 2.2 Flow aggregation
To compute an accurate flow map from noisy, empirical trajectories relies on resolving the crucial problem of how to aggregate similar, but not exactly overlapping, road segments. Many solutions have been offered, such as edge bundling (Zhou et al., 2013) and rasterisation (Wood et al., 2010, Morgan and Lovelace, 2021). Edge bundling consists of clustering trajectory linestrings and replacing all cluster members with a single representative linestring. These (and subsequent) authors conclude that it performs poorly when applied to noisy trajectories at the road level, and remains mostly suited to large-scale aggregations, such as desire line maps (Tobler, 1987). We do not consider edge bundling further since it is not suitable for our purposes. Rasterisation relies on converting the flow map into a raster matrix, and aggregating the flows within the same raster pixel neighbourhood. Whilst rasterisation is able to improve flow aggregation, it depends highly on the raster pixel neighbourhood size, and leads to a loss of spatial resolution.
We propose an alternative aggregation which solves both of these problems. We achieve this goal by local internal alignment between the map matched routes. By internal alignment, we mean that we align the routes with each other, rather than to the external road network graph. By local alignment, we mean that we align segments of the routes, rather than the complete routes. Our proposal is divided into several algorithms, which we develop within the R statistical analysis environment due to its integrated access to advanced statistical and geospatial analysis methods.
2.3 Node snapping and node splitting
@@
Since we are searching for points which are close to together, then this is well-suited to statistical clustering. There are many statistical clustering algorithms available, and we focus on hierarchical clustering. To resolve the computational bottleneck of hierarchical clustering, we approximate the 1-pass complete linkage clustering by a nested 2-pass clustering. We begin with an efficient single linkage clustering in the R package fastcluster (Müllner, 2013). Since single linkage can result in chain-like clusters, we compute a subsequent complete linkage clustering to break these potential chains. In this nested approach, the complete linkage distance matrix is calculated only within each single linkage cluster, and so it is less likely to reach computational limits.
Algorithm 2 is a description of ST_SNAPNODE. The inputs are the $n_{\mathcal{F}}$ flow linestrings =$\mathcal{F}=\{(f_{1},{\bm{f}}_{1}),\dots,(f_{n_{\mathcal{F}}},{\bm{f}}_{n_{\mathcal{F}}})\}$ and the snap tolerance $\varepsilon_{S}$. As its name suggests, all node points within $\varepsilon_{S}$ distance of each other are snapped together. In Step 1, we extract the boundary points of all flow linestrings. In Steps 2–3, we compute a single linkage clustering on all boundary points, and cut the dendrogram at height $\varepsilon_{S}$, resulting in $C^{\prime}$ clusters. In Steps 4–7, within each of these single linkage clusters, we compute a complete linkage clustering, and cut the dendrogram at height $\varepsilon_{S}$. This divides each single linkage cluster into $C^{\prime\prime}$ clusters, where all cluster members are at most $\varepsilon_{S}$ distance from each other. In Steps 8–11, we compute the weighted spatial centroid in each of the $C^{\prime\prime}$ complete linkage clusters, weighted by the traffic flow values. In Step 12, we renumber the cluster labels from the above nested clusterings to approximate a 1-pass complete linkage clustering into $C^{*}$ clusters. In Steps 13–16, within each of these $C^{*}$ clusters, for all points of the corresponding flow linestrings which are closer than $\varepsilon_{S}$ distance to the centroid, we snap them to the centroid. In Steps 17–18, we collate and sort all the $n_{\mathcal{F}^{*}}$ snapped linestrings =$\mathcal{F}^{*}=\{(f_{1}^{*},{\bm{f}}_{1}^{*}),\dots,(f_{n_{\mathcal{F}^{*}}}^{*},{\bm{f}}_{n_{\mathcal{F}^{*}}}^{*})\}$. In comparison to the unsnapped linestrings $\mathcal{F}$, there are fewer snapped linestrings $\mathcal{F}^{*}$ which have higher flow values, are more interconnected, and share more exactly overlapping sub-segments.
@@
3:Extract boundary points $B(\mathcal{F}):=\{\operatorname{Start}({\bm{f}}_{1}),\operatorname{End}({\bm{f}}_{1}),\dots,\operatorname{Start}({\bm{f}}_{n_{\mathcal{F}}}),\operatorname{End}({\bm{f}}_{n_{\mathcal{F}}})\}$
@@
16:     Extract corresponding flow linestrings $\mathcal{F}_{i}^{*}:=\{(f_{1}^{*},{\bm{f}}_{1}^{*}),\dots,(f_{n_{B^{*}}}^{*},{\bm{f}}_{n_{B^{*}}}^{*})\}$ from $\mathcal{F}$
17:     Snap points of $\{{\bm{f}}_{1}^{*},\dots,{\bm{f}}_{n_{B^{*}}}^{*}\}$ within $\varepsilon_{S}$ distance of cluster centroid to ${\bm{b}}^{*}_{i}$
@@
19:Collate snapped linestrings $\{\mathcal{F}_{1}^{*},\dots\mathcal{F}_{C^{\prime*}}^{*}\}$ into :=$\mathcal{F}^{*}:=\{(f_{1}^{*},{\bm{f}}_{1}^{*}),\dots,(f_{n_{\mathcal{F}^{*}}}^{*},{\bm{f}}_{n_{\mathcal{F}^{*}}}^{*})\}$
@@
Node snapping is focused near the boundary points of the flow linestrings, so it does not address misalignments far from the boundary points. Moreover there remain some linestrings whose intersection is not correctly computed by ST_OVERLINE_PLANR. The solution to both of these problems is the explicit addition of the missing intersection nodes to these linestrings.
ST_SPLITNODE is the collation of a couple of algorithms which explicitly add the nodes at the intersections in the interior of linestrings, and splits these linestrings into simple linestring segments at these added nodes. This gives the name to this method. The inputs to Algorithm 3 are the flow linestrings $\mathcal{F}$ and the node splitting type $S$. Our method deploys for $S=$ “subdivision” in Steps 1–3, to_spatial_subdivision in the sfnetworks package (van der Meer et al., 2023), and for $S=$ “unary” in Steps 4–5, geos_unary_union in the geos package (Dunnington and Pebesma, 2023). The first option tends to find fewer intersection nodes and is similar to the intersection computation performed by ST_OVERLINE_PLANR. The second option finds more missing intersection nodes, and leads to more comprehensive flow aggregation. We require both types of node splitting to compute a locally aligned flow map.
@@
6:else if $S==$ “unary” then
@@
In Figure 2(a) are the flow segments without any node splitting or node snapping. There are missing intersection nodes, and there are intersection nodes which are close to each other. In Figure 2(b) are the flow segments after node splitting ($S=$ “unary”) and then node snapping (with snap tolerance $\varepsilon_{S}=4$ m). The result is that there are fewer flow segments with higher flow values. Due to the combined action of ST_SPLITNODE and ST_SNAPNODE, Figure 2(b) is an improvement over Figure 2(a).
Refer to caption Refer to caption
Figure 2: Unsplit/unsnapped and node split/node snapped flow segments, with snap tolerance $\varepsilon_{S}=4$ m. (a) Unsplit and unsnapped. (b) Node split and node snapped.
@@ 2.4 Line blending to align similar linestrings with local reference
So far we have focused on improving the alignment of linestrings by resolving inconsistencies at their intersections. We now focus on aligning linestrings more generally. For this, we require a comparison relationship to establish an order of alignment of nearby linestrings. For two flow linestrings $(f_{1},{\bm{f}}_{1})$ and $(f_{2},{\bm{f}}_{2})$ from a flow map $\mathcal{F}$, we define that ${\bm{f}}_{2}$ is a candidate to be aligned to the reference ${\bm{f}}_{1}$ at threshold $\varepsilon\geq 0$ if
@@
The condition on the flow values $f_{2}\leq f_{1}$ means that we place a higher priority on the reference linestring ${\bm{f}}_{1}$ since its flow $f_{1}$ is larger than the flow $f_{2}$ on the candidate linestring ${\bm{f}}_{2}$. The buffer zone has flat edges, e.g. ST_BUFFER(endCapStyle="FLAT") in the R package sf (Pebesma, 2018), so the buffer zone ends at the boundary points of ${\bm{f}}_{1}$. This avoids erroneously considering neighbouring segments of ${\bm{f}}_{1}$, which are connected to ${\bm{f}}_{1}$ as a part of a longer sequence, to be candidates to be aligned to ${\bm{f}}_{1}$.
Let ${\bm{f}}_{\rm ref}$ be a reference linestring from $\mathcal{F}$. The set of $n_{\mathcal{F}_{\rm cand}}$ linestrings from \$\mathcal{F}\backslash\{{\bm{f}}_{\rm ref}\}$ which satisfy Equation (1) are the candidate linestrings $\mathcal{F}_{\rm cand}=\{{\bm{f}}_{{\rm cand},1},\dots,{\bm{f}}_{{\rm cand},n_{\mathcal{F}_{\rm cand}}}\}$, with the convention that if $n_{\mathcal{F}_{\rm cand}}=0$ then $\mathcal{F}_{\rm cand}$ is the empty set. We call our approach “line blending”, since we blend the candidate linestrings onto the reference linestring. The inputs in Algorithm 4 are the reference linestring ${\bm{f}}_{\rm ref}$, the $n_{\mathcal{F}_{\rm cand}}$ candidate linestrings $\mathcal{F}_{\rm cand}$, and the blend tolerance $\varepsilon$. The output are the modified reference linestring and $n_{\mathcal{F}_{\rm cand}}$ modified candidate flow linestrings, all with added interior points for accurate calculation of exactly equal linestring segments. In Step 1, we initialize a local network graph $\mathcal{N}^{*}$ with the reference linestring. In Steps 2–3, we extract all points of the candidate linestrings, and use the network_blend function in the sfnetworks package (which we denote as ST_NETWORK_BLEND) to blend efficiently these points into the reference linestring (van der Meer et al., 2023). ST_NETWORK_BLEND projects the candidate linestrings onto the reference linestring, and explicitly adds them to the network, thereby creating new edges in $\mathcal{N}^{*}$. The result is a local network graph with more, shorter edges and with nodes at the projected candidate points, and whose union is the reference linestring. In Steps 4–5, we extract the $n_{\mathcal{F}_{\rm cand}}$ blended candidate linestrings with these added interior points by applying the shortest path search between the start and end point of each blended candidate linestring along the network graph $\mathcal{N}^{*}$. Step 6 involves collating the blended candidate linestrings. Step 7 is the equivalent for the blended reference linestring.
@@
2:Output: ${\bm{f}}^{*}_{\rm ref}$ blended reference, $\mathcal{F}^{*}_{\rm cand}$ blended candidate flow linestrings
@@
8:Collate $\mathcal{F}^{*}_{\rm cand}:=\{{\bm{f}}_{{\rm cand},1}^{*},\dots,{\bm{f}}_{{\rm cand},n_{\mathcal{F}_{\rm cand}}}^{*}\}$
@@
9:${\bm{f}}^{*}_{{\rm ref}}$ := shortest path from $\operatorname{Start}({\bm{f}}_{{\rm ref}})$ to $\operatorname{End}({\bm{f}}_{{\rm ref}})$ along network $\mathcal{N}^{*}$
@@
The result from ${\tt ST\_LINEBLEND}$ is the modified reference linestring and the projected candidate linestrings with added interior points. If there are many candidate linestrings, then this may lead to many sub-segments in the projected linestrings, each with their own flow values. So we assign the rounded value of the weighted mean flow, weighted by the sub-segment length, to all sub-segments.
@@
In Figure 3(a), we illustrate this line blending with the flow linestrings $(7,ABC)$ in blue and $(2,AD)$ in orange. As these two linestrings do not share a common sub-segment, so flow aggregation does not modify them. Since $AD\subset{\tt ST\_BUFFER}(ABC,4\,\mathrm{m})$ (the pale blue rectangle), and $2<7$, then $ABC$ is the reference ${\bm{f}}_{\rm ref}$ and $AD$ is the candidate linestring ${\bm{f}}_{\rm cand}$, according to Equation (1). Line blending involves projecting $AD$ to $ABC$, so $D$ is projected to $D^{\prime}$ (which lies exactly on the reference linestring), and $B$ is added explicitly to this projected linestring. We also add $D^{\prime}$ to the reference linestring. The reference linestring becomes ${\bm{f}}^{*}_{\rm ref}=(7,ABD^{\prime}C)$, and the candidate ${\bm{f}}^{*}_{\rm cand}=(2,ABD^{\prime})$. Now ${\bm{f}}^{*}_{\rm ref},{\bm{f}}^{*}_{\rm cand}$ share the sub-segment $ABD^{\prime}$ exactly, and ST_OVERLINE_PLANR gives the aggregated flows as $(9,ABD^{\prime})$ and $(2,D^{\prime}C)$. We take the rounded value of the weighted mean of these two flow linestrings. The result is a single linestring $(9,ABD^{\prime}C)$, as shown in Figure 3(b).
Refer to caption Refer to caption
@@
Figure 3: Line blending leads to correct flow aggregation. (a) Before line blending and flow aggregation. Reference linestring $(7,ABC)$ in blue, candidate linestring $(2,AD)$ in orange. (b) After blending candidate into reference linestring, with tolerance $\varepsilon=4$ m, and flow aggregation. Blended linestring is $(9,ABD^{\prime}C)$.
After aggregation, the projected point $D^{\prime}$ in $ABD^{\prime}C$ appears to be redundant. Suppose that we have another linestring $(3,ABC)$ without $D^{\prime}$. Then ST_OVERLINE_PLANR does not aggregate $(9,ABD^{\prime}C)$ with $(3,ABC)$ even if $D^{\prime}$ is exactly coincident on the segment $BC$, since this algorithm does not explicitly add any missing intersection nodes prior to aggregation (Morgan and Lovelace, 2021). It outputs three linestrings $(12,AB),(9,BC),(3,BD^{\prime}C)$, rather than a single linestring $(12,ABD^{\prime}C)$. So we retain these intermediate points in the flow linestrings to assist in subsequent flow aggregation.
@@
More complex situations arise when other linestrings touch the candidate linestring, but are not candidates themselves for blending. Since line blending projects the candidates to the reference linestring, then we have to also project these other linestrings to avoid leaving a gap in the updated flow map. This procedure ST_SNAP_CAND_TOUCH is outlined in Algorithm 5. The inputs are the reference linestring ${\bm{f}}_{\rm ref}$, the $n_{\mathcal{F}_{\rm cand}}$ candidate linestrings $\mathcal{F}_{\rm cand}$, the $n_{\mathcal{F}_{\rm ct}}$ candidate-touching linestrings $\mathcal{F}_{\rm ct}$, and the snap tolerance $\varepsilon_{S}$. In Steps 1–8, we iterate over each candidate-touching linestring. In Steps 2–3, we compute the intersections between the candidate-touching linestring and the boundary points of the candidate linestrings, and the respective distances. In Steps 4–7, if this intersection point is within $\varepsilon_{S}$ distance to the boundary points of ${\bm{f}}_{\rm ref}$, then we snap the candidate-touching linestring to the closest boundary point. In Step 8, if the intersection point is not within $\varepsilon_{S}$ distance, then we snap it to the nearest point on ${\bm{f}}_{\rm ref}$. These snapping operations are similar to those in Steps 15–16 in ST_SNAPNODE in Algorithm 2, and ensure that we maintain connectivity between $\mathcal{F}_{\rm ct}$ and ${\bm{f}}_{\rm ref}$. In Step 9, we collate these snapped linestrings.
@@
Figure 4(a) is an illustration of ST_SNAP_CAND_TOUCH with the reference $(7,ABC)$ (blue), the candidate linestring $(2,AD)$ (orange), and the blending buffer zone with blend tolerance $\varepsilon=4$ m (light blue). The candidate-touching linestring ${\bm{f}}_{\rm ct}=(5,DE)$ (purple) touches the candidate $AD$ at $D$. The orange line is entirely within the pale blue buffer zone, whereas the purple line extends outside of the buffer zone and so is not a candidate for blending to the reference linestring. If we apply ST_LINEBLEND to blend the candidate linestring, then $D$ is projected to $D^{\prime}$ to lie on the reference linestring, i.e., ${\bm{f}}_{\rm cand}^{*}=(2,ABD^{\prime})$. If we apply ST_SNAP_CAND_TOUCH with snap tolerance $\varepsilon_{S}=4$ m to the candidate-touching linestring, then the intersection point $D$ is projected to the boundary point $C$, i.e., ${\bm{f}}^{*}_{\rm ct}=(5,CE)$. Thus line blending, candidate-touching snapping, and flow aggregation yields that ${\bm{f}}^{*}_{\rm ct}=(5,CE)$ remains connected to ${\bm{f}}_{\rm ref}^{*}=(9,ABD^{\prime}C)$.
Refer to caption Refer to caption